Graph each linear equation by finding and plotting its intercept points.
x-y=3
x-y=-4
x=5y
2x=y

Answer by MRperkins(300)

This is a rather simple problem once you understand what it is asking.
the y-intercept of a point is the y value when x=0
the x-intercept of a point is the x value when y=0
x-y=3
x-intercept:
when y=0 x=?
substitute 0 for y
x- y =3
x-(0)=3 or x=3
therefore the x-intercept of x-y=3 is the point (3,0)
y-intercept
when x=0 y=?
substitute 0 for x
x -y=3
(0)-y=3 or -y=3
divide both sides by -1 and you get y=-3
therefore the y-intercept of x-y=3 is the point (0,-3)
Graph these two points and then draw a straight line between them and this is your graph.
I beleive you will be able to successfully complete the other examples now. If you send me your answers to justin.sheppard.tech@hotmail.com I will check them for you.
.
.
Feedback is appreciated.
PayPal donations accepted through justin.sheppard.tech@hotmail.com